# A novel measurement of the strong-phase difference between $D^{0}\to K^{-}{{\pi}^{+}}$ and $\bar{D}^{0}\to K^{-}{{\pi}^{+}}$ decays using $C$-even and $C$-odd quantum-correlated ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pairs

### v1

@@
A novel measurement of the strong-phase difference between $D^{0}\to K^{-}{{\pi}^{+}}$ and $\bar{D}^{0}\to K^{-}{{\pi}^{+}}$ decays using $C$-even and $C$-odd quantum-correlated ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pairs
@@ I Introduction
and hence the decays of the two $D$ mesons are correlated. In ${{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}}\to Y_{1}\ \text{vs.}\ Y_{2}$ decays, where $Y_{1}$ and $Y_{2}$ are modes accessible to both the $D^{0}$ and $\bar{D}^{0}$ mesons, interference occurs between the two final state paths $D^{0}\to Y_{1},\bar{D}^{0}\to Y_{2}$ and $D^{0}\to Y_{2},\bar{D}^{0}\to Y_{1}$. The effects on the interference can be exploited to determine the hadronic parameters associated with the decay amplitudes of $Y_{1}$ and $Y_{2}$.
@@
Table 1: Mechanisms by which quantum-correlated ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pairs with eigenvalue $C$ are produced through the $e^{+}e^{-}\to{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;,\ D^{*}\bar{D}$ and $D^{*}\bar{D}^{*}$ processes.
@@
$e^{+}e^{-}\to\hskip 0.85358ptD^{*}\bar{D}\hskip 3.41432pt\to{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;{{\pi}^{0}}\phantom{\gamma*}$ $-1$
$e^{+}e^{-}\to D^{*}\bar{D}^{*}\to{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;\gamma\gamma\phantom{**}$ $-1$
$e^{+}e^{-}\to D^{*}\bar{D}^{*}\to{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;{{\pi}^{0}}\gamma\phantom{*}$ $+1$
$e^{+}e^{-}\to D^{*}\bar{D}^{*}\to{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;{{\pi}^{0}}{{\pi}^{0}}$ $-1$
The quantum-correlated ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pairs reconstructed in the dataset are exploited to measure the strong-phase difference between $D^{0}\to{{\textit{K}}^{-}}{{\pi}^{+}}$ and $\bar{D}^{0}\to{{\textit{K}}^{-}}{{\pi}^{+}}$ decays, denoted $\delta_{K\pi}^{D}$, and defined through the amplitude ratio,
${r_{K\pi}^{D}}e^{-i\delta_{K\pi}^{D}}=\frac{A(\bar{D}^{0}\to{{\textit{K}}^{-}}{{\pi}^{+}})}{A(D^{0}\to{{\textit{K}}^{-}}{{\pi}^{+}})},$ (3)
@@
$\bar{A}_{i}^{2}$ $=\int|A(\bar{D}^{0}\to Y_{i})|^{2}{\mathrm{d{\mathbf{p}}}},$ (6)
@@
$R_{i}e^{i\delta_{i}}$ $=\frac{\int A(D^{0}\to Y_{i})A(\bar{D}^{0}\to Y_{i})^{\ast}{\mathrm{d{\mathbf{p}}}}}{\sqrt{A_{i}^{2}\bar{A}_{i}^{2}}},$ (8)
@@ III Selection Requirements
Figure 1: The (a) $E_{\text{miss}}$, (b) $M_{\text{rec,}D}$, (c) $M_{\text{miss,${D}\kern 0.70004pt\overline{\kern-0.70004pt{D}}\;$}}^{2}$, (d) $M_{\text{miss,$\gamma{D}\kern 0.70004pt\overline{\kern-0.70004pt{D}}\;$}}^{2}$ distributions of ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;\to{{\textit{K}}^{-}}{{\pi}^{+}}\ \mathrm{vs.}\ {{\textit{K}}^{+}}{{\pi}^{-}}$ candidates in MC simulation samples. The $\Delta M_{{{D}^{*}}}$ vs. $M_{\text{rec,}{{D}^{*}}}$ distribution of simulated ${{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}}\to{{\textit{K}}^{-}}{{\pi}^{+}}\ \text{vs.}\ {{\textit{K}}^{+}}{{\pi}^{-}}$ decays which originate from the (e) $D^{*}\bar{D}^{*}\to\gamma\gamma{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ and $D^{*}\bar{D}^{*}\to\gamma{{\pi}^{0}}{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ and (f) $D^{*}\bar{D}^{*}\to{{\pi}^{0}}{{\pi}^{0}}{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ processes are also shown. The vertical dashed lines are used to distinguish between production mechanisms and correspond to the requirements in Table 2.
In each event the $D$-decay candidates are reconstructed using the above criteria. From these, a single best ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ candidate is chosen to be that with an average reconstructed invariant mass closest to the known $D^{0}$ mass [27], which is subsequently used to discriminate between the different production mechanisms. The strategy is first to isolate the $e^{+}e^{-}\to{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$, $e^{+}e^{-}\to D^{*}\bar{D}$ and $e^{+}e^{-}\to D^{*}\bar{D}^{*}$ processes, and then the possible ${{D}^{*}}$ decays.
@@
where $E_{\text{cm}}$ is the collision energy in the center-of-mass frame and $E_{{D}\kern 0.70004pt\overline{\kern-0.70004pt{D}}\;}$ is the energy of the ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pair calculated using the four-momentum, $p$, of each particle determined in an event-wide kinematic fit where all the hypothesized intermediate particles (${K^{0}_{\mathrm{S}}},{{\pi}^{0}},D^{0},D^{*0}$), are constrained to their known masses, $m$, from the PDG [27]. Figure 1(a) shows the $E_{\text{miss}}$ distribution of ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;\to{{\textit{K}}^{-}}{{\pi}^{+}}\ \mathrm{vs.}\ {{\textit{K}}^{+}}{{\pi}^{-}}$ candidates in simulation. This final state is used throughout the paper as a control channel due to the large branching fraction and reconstruction efficiency of the $D^{0}\to{{\textit{K}}^{-}}{{\pi}^{+}}$ decay. Furthermore, both decays in the $D^{0}\to{{\textit{K}}^{+}}{{\pi}^{-}},\bar{D}^{0}\to{{\textit{K}}^{-}}{{\pi}^{+}}$ path to the final state are doubly Cabibbo suppressed compared to the alternative, $D^{0}\to{{\textit{K}}^{-}}{{\pi}^{+}},\bar{D}^{0}\to{{\textit{K}}^{+}}{{\pi}^{-}}$, so the quantum correlations have little impact on the decay rate. In Fig. 1(a), the missing energy peaks at zero for candidates that originate from the $e^{+}e^{-}\to{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ process, but is larger for the $e^{+}e^{-}\to D^{*}\bar{D}$ and $e^{+}e^{-}\to D^{*}\bar{D}^{*}$ processes due to the missing light neutral particles from the ${D}^{*}$ decays.
@@
is used to distinguish between the $e^{+}e^{-}\to D^{*}\bar{D}$ and $e^{+}e^{-}\to D^{*}\bar{D}^{*}$ processes. The simulated $\Delta M_{\text{rec,}{D}}$ distribution for ${{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}}\to{{\textit{K}}^{-}}{{\pi}^{+}}\ \text{vs.}\ {{\textit{K}}^{+}}{{\pi}^{-}}$ candidates that pass the $E_{\text{miss}}>70\text{ MeV}$ requirement, to remove the contribution from the $e^{+}e^{-}\to{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ process, can be seen in Fig. 1(b). The recoil mass, $M_{\text{rec,}D}$, is the invariant mass of the other particles in the event after one of the $D$ meson decays is reconstructed with momentum $\vec{p}_{D}$,
@@
It is possible to use $E_{\text{miss}}$ to separate the $D^{*}\bar{D}$ processes from $D^{*}\bar{D}^{*}$, however, $M_{\text{rec,}{D}}$ displays a better resolution because it only relies on the kinematic information of a single $D$ meson from each event.
@@
Two sources of backgrounds pass the $E_{\text{miss}}$ and $M_{\text{rec,}{D}}$ requirements. Firstly, there are contributions from incorrectly reconstructed $D$ decays, which peak underneath the signal due to the $D$-meson mass constraints applied when calculating $E_{\text{miss}}$ and $M_{\text{rec,}{D}}$. These are subtracted from the samples through fits to the reconstructed invariant-masses of both $D$-decays as discussed in Sec. IV. Secondly, $D^{0}\bar{D}^{0}$ pairs which decay to the correct final state and are produced in the $e^{+}e^{-}\to{{D}^{*+}}{{D}^{*-}}$ process, where ${{D}^{*+}}\to{{\pi}^{+}}D^{0}$, pass the requirements to isolate the ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pairs from $D^{*}\bar{D}^{*}$ production with large rates. These are indistinguishable from signal in the aforementioned fits. However, they are significantly reduced by applying a requirement on the charged pion with the lowest momentum in each event, and through the criteria which isolate the production mechanisms that originate from the $D^{*}\bar{D}^{*}$ process discussed below.
The selection efficiency of the neutral particle from the ${D}^{*}$ decay is low. Therefore, to maximize the signal yields, the ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ production mechanisms that originate from the $D^{*}\bar{D}$ and $D^{*}\bar{D}^{*}$ processes are distinguished by partial reconstruction. For example, to discriminate between the $D^{*0}\to\gamma D^{0}$ and $D^{*0}\to{{\pi}^{0}}D^{0}$ decays that originate from the $e^{+}e^{-}\to D^{*}\bar{D}$ process, a missing mass variable, defined as
@@
Up to this point, separation between the ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$, $D^{*}\bar{D}\to{\gamma}{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}$ and $D^{*}\bar{D}\to{\pi^{0}}{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}$ production mechanisms has been achieved using three discriminating variables which only require reconstruction of the ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pair. However, the same strategy is much less effective in isolating the three production mechanisms that arise from $D^{*}\bar{D}^{*}$ decays due to the two additional neutral particles. Again, to maintain a high efficiency a partial reconstruction technique is employed. A single ${{D}^{*}}\to\gamma{D}$ candidate from each event, chosen to be that which has a reconstructed invariant-mass closest to the $D^{*0}$ from all combinations of the two selected ${D}$ mesons and all photon candidates, is used to distinguish the $D^{*}\bar{D}^{*}\to\gamma\gamma{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ and $D^{*}\bar{D}^{*}\to{{\pi}^{0}}\gamma{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ production mechanisms from $D^{*}\bar{D}^{*}\to{{\pi}^{0}}{{\pi}^{0}}{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$. This is achieved by selecting a region in the two-dimensional phase space defined by the difference between the reconstructed invariant mass of the ${D}^{*}$ candidate and the PDG value [27], denoted $\Delta M_{{{D}^{*}}}$, and the recoil mass of the ${D}^{*}$ candidate, $M_{\text{rec,}{{D}^{*}}}$, which is defined similarly to Eq. (11).
@@
For ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pairs produced in $D^{*}\bar{D}^{*}\to\gamma\gamma{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ and $D^{*}\bar{D}^{*}\to{{\pi}^{0}}\gamma{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ decays, the corresponding distributions will peak around $\Delta M_{{{D}^{*}}}\simeq 0\text{ GeV}\!/c^{2}$ and $M_{\text{rec,}{{D}^{*}}}\simeq m_{{D}^{*}}$, as shown in Fig. 1(e). The long tail at $\Delta M_{{{D}^{*}}}\simeq 0$ and high $M_{\text{rec,}{{D}^{*}}}$ occurs due to ISR, whilst the other, shorter, tails are caused by the detector resolution. For the $D^{*}\bar{D}^{*}\to{{\pi}^{0}}{{\pi}^{0}}{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ production mechanism, the ${D}^{*}$ meson is reconstructed by combining a $D$ meson with a random photon, and thus the corresponding distribution will peak in a displaced region dictated by the momentum distribution of the photon, as can be seen in Fig. 1(f). Due to the correlation of the $\Delta M_{{{D}^{*}}}$ and $M_{\text{rec,}{{D}^{*}}}$ variables for $D^{*}\bar{D}^{*}\to{{\pi}^{0}}{{\pi}^{0}}{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ events, non-rectangular selections are applied to remove misidentified $D^{*}\bar{D}^{*}\to{{\pi}^{0}}{{\pi}^{0}}{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ events. The $D^{0}\bar{D}^{0}$ pairs produced in the ${{D}^{*+}}{{D}^{*-}}$ process show the same displacement, and therefore they are effectively removed by the same selections from the $\gamma\gamma{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ and $\gamma{{\pi}^{0}}{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ samples.
@@
The $D^{*}\bar{D}^{*}\to\gamma\gamma{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ production mechanism is distinguished from $D^{*}\bar{D}^{*}\to{{\pi}^{0}}\gamma{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ using the missing mass after reconstructing the ${D}$ and $\bar{{D}}^{*}$ mesons,

### v2

# A novel measurement of the strong-phase difference between ${{D}^{0}}\to K^{-}{{\pi}^{+}}$ and ${{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{0}}\to K^{-}{{\pi}^{+}}$ decays using $C$-even and $C$-odd quantum-correlated ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pairs

@@
A novel measurement of the strong-phase difference between ${{D}^{0}}\to K^{-}{{\pi}^{+}}$ and ${{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{0}}\to K^{-}{{\pi}^{+}}$ decays using $C$-even and $C$-odd quantum-correlated ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pairs
@@ I Introduction
and hence the decays of the two $D$ mesons are correlated. In ${{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}}\to Y_{1}\ \text{vs.}\ Y_{2}$ decays, where $Y_{1}$ and $Y_{2}$ are modes accessible to both the ${D}^{0}$ and ${\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{0}$ mesons, interference occurs between the two final state paths ${{D}^{0}}\to Y_{1},{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{0}}\to Y_{2}$ and ${{D}^{0}}\to Y_{2},{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{0}}\to Y_{1}$. The effects on the interference can be exploited to determine the hadronic parameters associated with the decay amplitudes of $Y_{1}$ and $Y_{2}$.
@@
Table 1: Mechanisms by which quantum-correlated ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pairs with eigenvalue $C$ are produced through the $e^{+}e^{-}\to{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;,\ {{D}^{*}}\kern 1.00006pt\overline{\kern-1.00006pt{D}}$ and ${{D}^{*}}{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{*}}$ processes.
@@
$e^{+}e^{-}\to\hskip 0.85358pt{{D}^{*}}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\hskip 3.41432pt\to{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;{{\pi}^{0}}\phantom{\gamma*}$ $-1$
$e^{+}e^{-}\to{{D}^{*}}{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{*}}\to{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;\gamma\gamma\phantom{**}$ $-1$
$e^{+}e^{-}\to{{D}^{*}}{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{*}}\to{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;{{\pi}^{0}}\gamma\phantom{*}$ $+1$
$e^{+}e^{-}\to{{D}^{*}}{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{*}}\to{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;{{\pi}^{0}}{{\pi}^{0}}$ $-1$
The quantum-correlated ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pairs reconstructed in the dataset are exploited to measure the strong-phase difference between ${{D}^{0}}\to{{\textit{K}}^{-}}{{\pi}^{+}}$ and ${{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{0}}\to{{\textit{K}}^{-}}{{\pi}^{+}}$ decays, denoted $\delta_{K\pi}^{D}$, and defined through the amplitude ratio,
${r_{K\pi}^{D}}e^{-i\delta_{K\pi}^{D}}=\frac{A({{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{0}}\to{{\textit{K}}^{-}}{{\pi}^{+}})}{A({{D}^{0}}\to{{\textit{K}}^{-}}{{\pi}^{+}})},$ (3)
@@
$\bar{A}_{i}^{2}$ $=\int|A({{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{0}}\to Y_{i})|^{2}{\mathrm{d{\mathbf{p}}}},$ (6)
@@
$R_{i}e^{i\delta_{i}}$ $=\frac{\int A({{D}^{0}}\to Y_{i})A({{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{0}}\to Y_{i})^{\ast}{\mathrm{d{\mathbf{p}}}}}{\sqrt{A_{i}^{2}\bar{A}_{i}^{2}}},$ (8)
@@ III Selection Requirements
Figure 1: The (a) $E_{\text{miss}}$, (b) $M_{\text{rec,}D}$, (c) $M_{\text{miss,${D}\kern 0.70004pt\overline{\kern-0.70004pt{D}}\;$}}^{2}$, (d) $M_{\text{miss,$\gamma{D}\kern 0.70004pt\overline{\kern-0.70004pt{D}}\;$}}^{2}$ distributions of ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;\to{{\textit{K}}^{-}}{{\pi}^{+}}\ \mathrm{vs.}\ {{\textit{K}}^{+}}{{\pi}^{-}}$ candidates in MC simulation samples. The $\Delta M_{{{D}^{*}}}$ vs. $M_{\text{rec,}{{D}^{*}}}$ distribution of simulated ${{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}}\to{{\textit{K}}^{-}}{{\pi}^{+}}\ \text{vs.}\ {{\textit{K}}^{+}}{{\pi}^{-}}$ decays which originate from the (e) ${{D}^{*}}{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{*}}\to\gamma\gamma{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ and ${{D}^{*}}{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{*}}\to\gamma{{\pi}^{0}}{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ and (f) ${{D}^{*}}{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{*}}\to{{\pi}^{0}}{{\pi}^{0}}{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ processes are also shown. The vertical dashed lines are used to distinguish between production mechanisms and correspond to the requirements in Table 2.
In each event the $D$-decay candidates are reconstructed using the above criteria. From these, a single best ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ candidate is chosen to be that with an average reconstructed invariant mass closest to the known ${{D}^{0}}$ mass [27], which is subsequently used to discriminate between the different production mechanisms. The strategy is first to isolate the $e^{+}e^{-}\to{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$, $e^{+}e^{-}\to{{D}^{*}}\kern 1.00006pt\overline{\kern-1.00006pt{D}}$ and $e^{+}e^{-}\to{{D}^{*}}{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{*}}$ processes, and then the possible ${{D}^{*}}$ decays.
@@
where $E_{\text{cm}}$ is the collision energy in the center-of-mass frame and $E_{{D}\kern 0.70004pt\overline{\kern-0.70004pt{D}}\;}$ is the energy of the ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pair calculated using the four-momentum, $p$, of each particle determined in an event-wide kinematic fit where all the hypothesized intermediate particles (${K^{0}_{\mathrm{S}}},{{\pi}^{0}},{{D}^{0}},D^{*0}$), are constrained to their known masses, $m$, from the PDG [27]. Figure 1(a) shows the $E_{\text{miss}}$ distribution of ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;\to{{\textit{K}}^{-}}{{\pi}^{+}}\ \mathrm{vs.}\ {{\textit{K}}^{+}}{{\pi}^{-}}$ candidates in simulation. This final state is used throughout the paper as a control channel due to the large branching fraction and reconstruction efficiency of the ${{D}^{0}}\to{{\textit{K}}^{-}}{{\pi}^{+}}$ decay. Furthermore, both decays in the ${{D}^{0}}\to{{\textit{K}}^{+}}{{\pi}^{-}},{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{0}}\to{{\textit{K}}^{-}}{{\pi}^{+}}$ path to the final state are doubly Cabibbo suppressed compared to the alternative, ${{D}^{0}}\to{{\textit{K}}^{-}}{{\pi}^{+}},{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{0}}\to{{\textit{K}}^{+}}{{\pi}^{-}}$, so the quantum correlations have little impact on the decay rate. In Fig. 1(a), the missing energy peaks at zero for candidates that originate from the $e^{+}e^{-}\to{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ process, but is larger for the $e^{+}e^{-}\to{{D}^{*}}\kern 1.00006pt\overline{\kern-1.00006pt{D}}$ and $e^{+}e^{-}\to{{D}^{*}}{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{*}}$ processes due to the missing light neutral particles from the ${D}^{*}$ decays.
@@
is used to distinguish between the $e^{+}e^{-}\to{{D}^{*}}\kern 1.00006pt\overline{\kern-1.00006pt{D}}$ and $e^{+}e^{-}\to{{D}^{*}}{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{*}}$ processes. The simulated $\Delta M_{\text{rec,}{D}}$ distribution for ${{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}}\to{{\textit{K}}^{-}}{{\pi}^{+}}\ \text{vs.}\ {{\textit{K}}^{+}}{{\pi}^{-}}$ candidates that pass the $E_{\text{miss}}>70\text{ MeV}$ requirement, to remove the contribution from the $e^{+}e^{-}\to{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ process, can be seen in Fig. 1(b). The recoil mass, $M_{\text{rec,}D}$, is the invariant mass of the other particles in the event after one of the $D$ meson decays is reconstructed with momentum $\vec{p}_{D}$,
@@
It is possible to use $E_{\text{miss}}$ to separate the ${{D}^{*}}\kern 1.00006pt\overline{\kern-1.00006pt{D}}$ processes from ${{D}^{*}}{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{*}}$, however, $M_{\text{rec,}{D}}$ displays a better resolution because it only relies on the kinematic information of a single $D$ meson from each event.
@@
Two sources of backgrounds pass the $E_{\text{miss}}$ and $M_{\text{rec,}{D}}$ requirements. Firstly, there are contributions from incorrectly reconstructed $D$ decays, which peak underneath the signal due to the $D$-meson mass constraints applied when calculating $E_{\text{miss}}$ and $M_{\text{rec,}{D}}$. These are subtracted from the samples through fits to the reconstructed invariant-masses of both $D$-decays as discussed in Sec. IV. Secondly, ${{D}^{0}}{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{0}}$ pairs which decay to the correct final state and are produced in the $e^{+}e^{-}\to{{D}^{*+}}{{D}^{*-}}$ process, where ${{D}^{*+}}\to{{\pi}^{+}}{{D}^{0}}$, pass the requirements to isolate the ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pairs from ${{D}^{*}}{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{*}}$ production with large rates. These are indistinguishable from signal in the aforementioned fits. However, they are significantly reduced by applying a requirement on the charged pion with the lowest momentum in each event, and through the criteria which isolate the production mechanisms that originate from the ${{D}^{*}}{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{*}}$ process discussed below.
The selection efficiency of the neutral particle from the ${D}^{*}$ decay is low. Therefore, to maximize the signal yields, the ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ production mechanisms that originate from the ${{D}^{*}}\kern 1.00006pt\overline{\kern-1.00006pt{D}}$ and ${{D}^{*}}{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{*}}$ processes are distinguished by partial reconstruction. For example, to discriminate between the $D^{*0}\to\gamma{{D}^{0}}$ and $D^{*0}\to{{\pi}^{0}}{{D}^{0}}$ decays that originate from the $e^{+}e^{-}\to{{D}^{*}}\kern 1.00006pt\overline{\kern-1.00006pt{D}}$ process, a missing mass variable, defined as
@@
Up to this point, separation between the ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$, $D^{*}\bar{D}\to{\gamma}{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}$ and $D^{*}\bar{D}\to{\pi^{0}}{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}$ production mechanisms has been achieved using three discriminating variables which only require reconstruction of the ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pair. However, the same strategy is much less effective in isolating the three production mechanisms that arise from ${{D}^{*}}{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{*}}$ decays due to the two additional neutral particles. Again, to maintain a high efficiency a partial reconstruction technique is employed. A single ${{D}^{*}}\to\gamma{D}$ candidate from each event, chosen to be that which has a reconstructed invariant-mass closest to the $D^{*0}$ from all combinations of the two selected ${D}$ mesons and all photon candidates, is used to distinguish the ${{D}^{*}}{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{*}}\to\gamma\gamma{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ and ${{D}^{*}}{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{*}}\to{{\pi}^{0}}\gamma{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ production mechanisms from ${{D}^{*}}{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{*}}\to{{\pi}^{0}}{{\pi}^{0}}{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$. This is achieved by selecting a region in the two-dimensional phase space defined by the difference between the reconstructed invariant mass of the ${D}^{*}$ candidate and the PDG value [27], denoted $\Delta M_{{{D}^{*}}}$, and the recoil mass of the ${D}^{*}$ candidate, $M_{\text{rec,}{{D}^{*}}}$, which is defined similarly to Eq. (11).
@@
For ${D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ pairs produced in ${{D}^{*}}{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{*}}\to\gamma\gamma{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ and ${{D}^{*}}{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{*}}\to{{\pi}^{0}}\gamma{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ decays, the corresponding distributions will peak around $\Delta M_{{{D}^{*}}}\simeq 0\text{ GeV}\!/c^{2}$ and $M_{\text{rec,}{{D}^{*}}}\simeq m_{{D}^{*}}$, as shown in Fig. 1(e). The long tail at $\Delta M_{{{D}^{*}}}\simeq 0$ and high $M_{\text{rec,}{{D}^{*}}}$ occurs due to ISR, whilst the other, shorter, tails are caused by the detector resolution. For the ${{D}^{*}}{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{*}}\to{{\pi}^{0}}{{\pi}^{0}}{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ production mechanism, the ${D}^{*}$ meson is reconstructed by combining a $D$ meson with a random photon, and thus the corresponding distribution will peak in a displaced region dictated by the momentum distribution of the photon, as can be seen in Fig. 1(f). Due to the correlation of the $\Delta M_{{{D}^{*}}}$ and $M_{\text{rec,}{{D}^{*}}}$ variables for ${{D}^{*}}{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{*}}\to{{\pi}^{0}}{{\pi}^{0}}{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ events, non-rectangular selections are applied to remove misidentified ${{D}^{*}}{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{*}}\to{{\pi}^{0}}{{\pi}^{0}}{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ events. The ${{D}^{0}}{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{0}}$ pairs produced in the ${{D}^{*+}}{{D}^{*-}}$ process show the same displacement, and therefore they are effectively removed by the same selections from the $\gamma\gamma{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ and $\gamma{{\pi}^{0}}{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ samples.
@@
The ${{D}^{*}}{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{*}}\to\gamma\gamma{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ production mechanism is distinguished from ${{D}^{*}}{{\kern 1.00006pt\overline{\kern-1.00006ptD}}{}^{*}}\to{{\pi}^{0}}\gamma{D}\kern 1.00006pt\overline{\kern-1.00006pt{D}}\;$ using the missing mass after reconstructing the ${D}$ and $\bar{{D}}^{*}$ mesons,